# Turbulent cascade arrests and the formation of intermediate-scale condensates

### v1

@@ Introduction.
Since friction is absent, the forcing term is balanced by the transfer term at $k_{n}=k_{f}$, and, in the cascade range $k_{n}<k_{\mathrm{tr}}$, the contribution from the transfer term is negligible, so the statistical properties are like those in the pure 3D direct cascade ($k_{\mathrm{tr}}/k_{f}\to\infty$). Deviations from this 3D cascade arise when we account for the dissipation term. The maximum of $|D_{\nu}(k_{n})|$ shifts from $k_{n}\simeq k_{\nu}$ to $k_{n}\simeq k_{\mathrm{tr}}$, thus compensating for the accumulation of energy at the same wavenumbers, as we show in the inset of Fig. 2(c), where $T(k_{n})$ is balanced by $D_{\nu}(k_{n})$ at $k_{n}\simeq k_{\mathrm{tr}}$. Moreover, as $k_{\mathrm{tr}}$ decreases, the maximum of $|D_{\nu}(k_{n})|$ shifts to smaller values of $k_{n}$ and its magnitude increases to compensate for the stronger build-up of energy [Fig. 2(c)].
@@
$\Pi_{H}(k_{n})=\bigg{\langle}\sum^{N}_{j=n}\Re\big{\{}k_{n}^{\beta}\Phi_{j}u^{*}_{j}\big{\}}\bigg{\rangle},$ (8)
where $\beta=\log_{\lambda}(a/c)$. Clearly, as $k_{\mathrm{tr}}$ is decreased, the flux increases and tends to flatten for $k_{\mathrm{tr}}\ll k_{n}\ll k_{\nu}$, suggesting a direct cascade of $H$. However, the lack of significant separation between $k_{\mathrm{tr}}$ and $k_{\nu}$ makes it difficult to identify a range where $\Pi_{H}(k_{n})$ remains constant. In the inset of Fig. 3(a), we plot $\Pi_{H}(k_{n})$ for fixed $k_{\mathrm{tr}}$ and different values of $\nu$ to observe indeed that, for small viscosities, $\Pi_{H}(k_{n})$ tends to flatten for $k_{\mathrm{tr}}\ll k_{n}\ll k_{\nu}$.
As further confirmation of the direct cascade of $H$, we show in Fig. 3(b) that, by moving $k_{\mathrm{tr}}$ close to $k_{1}$, we achieve a large range of constant $\Pi_{H}(k_{n})$, in which $\mathcal{E}(k_{n})\sim k_{n}^{-\gamma}$ with $\gamma=2[1+\beta]/3+1$ [inset of Fig. 3(b)], as is expected in the direct-cascade regime of $H$ Gilbert et al. (2002).
@@
We now address the case $k_{f}>k_{\mathrm{tr}}$ [Table 1, runs D; $\mu\neq 0$ helps the system to reach a statistically stationary state]. For $k_{\mathrm{tr}}=0$ and with our choice of parameters, $\mathcal{E}(k_{n})\sim k_{n}^{-5/3},$ in the range between the friction-dominated wavenumbers and $k_{f}$ Gilbert et al. (2002). In this range, $\Pi_{E}(k_{n})<0$ remains constant and equals the rate of hyper-friction energy dissipation $\epsilon_{\mu}=\sum^{N}_{n=1}D_{\mu}(k_{n})$.
We now consider $k_{f}>k_{\mathrm{tr}}>0$. Small values of $k_{\mathrm{tr}}$ have a negligible effects on energy spectra and fluxes, because at small $k_{n}$ the inverse energy cascade is already stopped by hyper-friction; so we focus on intermediate values of $k_{\mathrm{tr}}$ between $k_{1}$ and $k_{f}$, where the energy flux [Fig. 4(a)] indicates that the inverse energy cascade is arrested at $k_{n}\simeq k_{\mathrm{tr}}$ and it is accompanied by energy build-up around $k_{\mathrm{tr}}$ [Fig. 4(b)].
We see the following two scaling forms on each side of $k_{\mathrm{tr}}$: For $k_{1}\ll k_{n}\ll k_{tr}$, $\mathcal{E}(k_{n})\sim k_{n}^{-1}$, which indicates equipartition [Fig. 4(b)]; and for $k_{\mathrm{tr}}\ll k_{n}\ll k_{f}$, $\mathcal{E}(k_{n})\sim k_{n}^{-5/3}$, as we expect in the range of the inverse energy cascade [inset of Fig. 4(b)]. Clearly, the latter range decreases as $k_{\mathrm{tr}}$ approaches $k_{f}$. The energy transfer that leads to a statistically stationary state is similar to that observed in case $(a)$: the accumulation of energy at scales comparable to $k_{\mathrm{tr}}$ is compensated by an increased viscous dissipation at similar scales. Indeed, in Fig. 4(a) we see, together with a peak of dissipation at $k_{f}$, a second peak at $k_{\mathrm{tr}}$  [^1] We have checked that the contribution of hyper-friction to the formation of the energy condensate is negligible..
@@
Figure 4: Plots for runs D1-D5: (a) Log-log plot of the compensated spectra $k_{n}\mathcal{E}(k_{n})$ versus $k_{n}/k_{\mathrm{tr}}$ and in the inset the compensated spectra $k_{n}^{5/3}\mathcal{E}(k_{n})$ versus $k_{n}/k_{\mathrm{tr}}$. (b) Log-linear plots of $\Pi(k_{n})$ versus $k_{n}/k_{\nu}$ and in the inset $D_{\nu}(k_{n})$ versus $k_{n}/k_{\nu}$.
Our study sheds new light on a general energy-transfer mechanism for the non-dissipative arrest of energy cascades (inverse or direct) when 3D turbulent dynamics, at small $k$, coexists with 2D turbulent dynamics, at large $k$. The shell-model approach we employ allows us to cover a large range of wavenumbers; this is crucial for uncovering the subtle interplay between the nonlinear and viscous terms. Specifically, we find that, when $k_{\mathrm{tr}}/k_{f}>1$, the direct energy cascade for $k_{n}<k_{\mathrm{tr}}$ is arrested by the 2D-like dynamics at $k_{n}>k_{\mathrm{tr}}$. In contrast, when $k_{\mathrm{tr}}/k_{f}<1$, the inverse energy cascade for $k_{n}>k_{\mathrm{tr}}$ is arrested by the 3D-like dynamics at $k_{n}<k_{\mathrm{tr}}$. In both cases, the arrest of the cascade, close to $k_{\mathrm{tr}}$, results in energy accumulation around $k_{\mathrm{tr}}$. In a spatially extended system, such an accumulation of energy would lead to the emergence of spatial structures, or condensates, of size $k_{\mathrm{tr}}^{-1}$. A statistically stationary state stems from an increased viscous dissipation, at wavenumbers close to $k_{\mathrm{tr}}$, which compensates for the energy accumulation. This is reminiscent of condensates in 2D turbulence. Furthermore, we show that, when $k_{\mathrm{tr}}/k_{f}>1$, the energy that accumulates near to $k_{\mathrm{tr}}$ generates a direct cascade of generalized enstrophy for $k_{n}>k_{\mathrm{tr}}$, whereas, when $k_{\mathrm{tr}}/k_{f}<1$, the modes with $k_{n}<k_{\mathrm{tr}}$ are in statistical equilibrium.
Our modifications to the shell model with scale-dependent coefficients is not restricted to a given physical system  [^2] A shell-model version of the 3D NSE with odd viscosity de Wit et al. (2024a) could be constructed by adding a term proportional to $-\mathrm{i}k_{n}^{2}u_{n}$ to the SABRA model, in the same spirit as shell models of rotating turbulence Rathor et al. (2022).. Although the details of the dependence of the transitional scale may vary with the parameters of specific models, the energy-transfer mechanisms that we have identified are general, so they should appear in any turbulent system that has a 2D-like large-$k$ dynamics and 3D-like small-$k$ dynamics. Thus, our results will stimulate the study of new physical systems with such cascade arrests that lead to intermediate-scale condensates.
@@ Acknowledgements.
[…]

### v2

@@ Introduction.
Here, $T(k_{n})=\langle\Re\{\mathrm{i}\Phi_{n}u_{n}^{*}\}\rangle,~{}~{}F(k_{n})=\langle\Re\{f_{n}u_{n}^{*}\}\rangle,~{}~{}D_{\nu}(k_{n})=\langle 2\nu k_{n}^{2}|u_{n}|^{2}\rangle,~{}{\rm{and}}~{}D_{\mu}(k_{n})=\langle 2\mu k_{n}^{-2}|u_{n}|^{2}\rangle$ are the nonlinear, forcing, viscous, and friction contributions, respectively. We plot these quantities in the inset of Fig. 2(c) for $n_{\mathrm{tr}}=10$. Since friction is absent, the forcing term is balanced by the transfer term at $k_{n}=k_{f}$, and, in the cascade range $k_{n}<k_{\mathrm{tr}}$, the contribution from the transfer term is negligible, so the statistical properties are like those in the pure 3D direct cascade ($k_{\mathrm{tr}}/k_{f}\to\infty$). Deviations from this 3D cascade arise when we account for the dissipation term. The maximum of $|D_{\nu}(k_{n})|$ shifts from $k_{n}\simeq k_{\nu}$ to $k_{n}\simeq k_{\mathrm{tr}}$, thus compensating for the accumulation of energy at the same wavenumbers. This is clearly shown in the inset of Fig. 2(c), where $T(k_{n})$ is balanced by $D_{\nu}(k_{n})$ at $k_{n}\simeq k_{\mathrm{tr}}$. Moreover, as $k_{\mathrm{tr}}$ decreases, the maximum of $|D_{\nu}(k_{n})|$ shifts to smaller values of $k_{n}$ and its magnitude increases to compensate for the stronger build-up of energy [Fig. 2(c)].
@@
$\Pi_{H}(k_{n})=\bigg{\langle}\sum^{N}_{j=n}\Re\big{\{}\mathrm{i}k_{n}^{\beta}\Phi_{j}u^{*}_{j}\big{\}}\bigg{\rangle}$ (8)
with $\beta=\log_{\lambda}(a/c)$. Clearly, as $k_{\mathrm{tr}}$ is decreased, the flux increases and tends to flatten for $k_{\mathrm{tr}}\ll k_{n}\ll k_{\nu}$, suggesting a direct cascade of $H$. However, the lack of significant separation between $k_{\mathrm{tr}}$ and $k_{\nu}$ makes it difficult to identify a range where $\Pi_{H}(k_{n})$ remains constant. In the inset of Fig. 3(a), we plot $\Pi_{H}(k_{n})$ for fixed $k_{\mathrm{tr}}$ and different values of $\nu$ to observe indeed that, for small viscosities, $\Pi_{H}(k_{n})$ tends to flatten for $k_{\mathrm{tr}}\ll k_{n}\ll k_{\nu}$. As further confirmation of the direct cascade of $H$, we show in Fig. 3(b) that, by moving $k_{\mathrm{tr}}$ close to $k_{1}$, we achieve a large range of constant $\Pi_{H}(k_{n})$. Moreover, in such range, $\mathcal{E}(k_{n})\sim k_{n}^{-\gamma}$ with $\gamma=2[1+\beta]/3+1$ [inset of Fig. 3(b)], as is expected in the direct-cascade regime of $H$ [29].
@@
We now address the case $k_{f}>k_{\mathrm{tr}}$ (Table 1, runs D; $\mu\neq 0$ helps the system to reach a statistically stationary state but does not contribute to the formation of the condensate  [^1] We have repeated runs D2-D5 with $\mu=0$, and the conclusions remain the same). For $k_{\mathrm{tr}}=0$ and with our choice of parameters, $\mathcal{E}(k_{n})\sim k_{n}^{-5/3},$ in the range between the friction-dominated wavenumbers and $k_{f}$ [29]. In this range, $\Pi_{E}(k_{n})<0$ remains constant and equals the rate of hyper-friction energy dissipation $\epsilon_{\mu}=\sum^{N}_{n=1}D_{\mu}(k_{n})$.
@@
We now consider $k_{f}>k_{\mathrm{tr}}>0$. Small values of $k_{\mathrm{tr}}$ have a negligible effect on energy spectra and fluxes, because at small $k_{n}$ the inverse energy cascade is already stopped by hyper-friction; so we focus on intermediate values of $k_{\mathrm{tr}}$ between $k_{1}$ and $k_{f}$. For such values of $k_{\mathrm{tr}}$, the energy flux [Fig. 4(a)] indicates that the inverse energy cascade is arrested at $k_{n}\simeq k_{\mathrm{tr}}$ and it is accompanied by energy build-up (i.e. the formation of an intermediate-scale condensate) around $k_{\mathrm{tr}}$ [Fig. 4(b)].
We see the following two scaling forms on each side of $k_{\mathrm{tr}}$: For $k_{1}\ll k_{n}\ll k_{tr}$, $\mathcal{E}(k_{n})\sim k_{n}^{-1}$, which indicates equipartition [Fig. 4(b)]; and for $k_{\mathrm{tr}}\ll k_{n}\ll k_{f}$, $\mathcal{E}(k_{n})\sim k_{n}^{-5/3}$, as we expect in the range of the inverse energy cascade [inset of Fig. 4(b)]. Clearly, the latter range decreases as $k_{\mathrm{tr}}$ approaches $k_{f}$. The energy transfer that leads to a statistically stationary state is similar to that observed in case $(a)$: the accumulation of energy at scales comparable to $k_{\mathrm{tr}}$ is compensated by an increased viscous dissipation at similar scales. Indeed, in Fig. 4(a) we see, together with a peak of dissipation at $k_{f}$, a second peak at $k_{\mathrm{tr}}$  [^2] We have checked that the contribution of hyper-friction to the formation of the energy condensate is negligible.. Before concluding, we comment on the aforementioned regime of energy equipartition. In shell models, with constant 3D-like coefficients, the wavenumbers $k_{n}\ll k_{f}$ are isolated from the effects of forcing and viscous damping, and the energy flux to these small wavenumbers is zero [34]. Therefore, energy in these modes is expected to equilibrate [1, 34]. Analogous arguments lead to energy equipartition for $k_{n}\ll k_{\mathrm{tr}}$ in our model: the condensate generated around $k_{\mathrm{tr}}$ by the large-$k_{n}$ inverse cascade indeed acts as a source of energy for the 3D-like low-$k_{n}$ modes. It remains to be understood whether or not such a regime is expected to persist in real turbulent flows. Energy equipartition at length scales larger than the forcing length scale has been reported in both numerical and experimental investigations [35, 36, 37]. However, a recent study [38] has shown that, in 3D homogeneous and isotropic turbulence, the statistics of the large scales is not Gaussian. This has been attributed to a strong coupling between the small and large scales, which may not be captured by shell models. Furthermore, numerical simulations of 3D turbulence have reported significant deviations from the equilibrium spectrum for certain forcing choices [36, 39]. Therefore, we expect that spatial and temporal features of the intermediate-scale condensate, specific to the hydrodynamical PDEs, are detrimental to the presence of equipartition at large length scales.
@@
Figure 4: Plots for runs D1-D5: (a) Log-log plot of the compensated spectra $k_{n}\mathcal{E}(k_{n})$ versus $k_{n}/k_{\mathrm{tr}}$ and in the inset the compensated spectra $k_{n}^{5/3}\mathcal{E}(k_{n})$ versus $k_{n}/k_{\mathrm{tr}}$. (b) Log-linear plots of $\Pi(k_{n})$ versus $k_{n}/k_{\nu}$ and in the inset $D_{\nu}(k_{n})$ versus $k_{n}/k_{\nu}$. The Kolmogorov wavenumber corresponds to $n$ lying between 27 and 28
Conclusions.
Our study identifies a general energy-transfer mechanism for the non-dissipative arrest of energy cascades (inverse or direct) when 3D turbulent dynamics, at small $k$, coexists with 2D turbulent dynamics, at large $k$. The shell-model approach we employ allows us to resolve a large range of wavenumbers; this is crucial for uncovering the subtle interplay between the nonlinear and viscous terms. Specifically, we find that, when $k_{\mathrm{tr}}/k_{f}>1$, the direct energy cascade for $k_{n}<k_{\mathrm{tr}}$ is arrested by the 2D-like dynamics at $k_{n}>k_{\mathrm{tr}}$. In contrast, when $k_{\mathrm{tr}}/k_{f}<1$, the inverse energy cascade for $k_{n}>k_{\mathrm{tr}}$ is arrested by the 3D-like dynamics at $k_{n}<k_{\mathrm{tr}}$. In both cases, the arrest close to $k_{\mathrm{tr}}$, results in energy accumulation around $k_{\mathrm{tr}}$. In a spatially extended system, such an accumulation of energy would lead to the emergence of spatial structures, or condensates, of size $k_{\mathrm{tr}}^{-1}$. A statistically stationary state stems from an increased viscous dissipation, at wavenumbers close to $k_{\mathrm{tr}}$, which compensates for the energy accumulation. This is reminiscent of condensates in 2D turbulence. Furthermore, we show that, when $k_{\mathrm{tr}}/k_{f}>1$, the energy that accumulates near to $k_{\mathrm{tr}}$ generates a direct cascade of generalized enstrophy for $k_{n}>k_{\mathrm{tr}}$, whereas, when $k_{\mathrm{tr}}/k_{f}<1$, the modes with $k_{n}<k_{\mathrm{tr}}$ are in statistical equilibrium.
The details of the transitional scale will vary with different models, but the energy-transfer mechanisms we have identified are general and apply to any turbulent system with 2D-like large-$k$ dynamics and 3D-like small-$k$ dynamics. Thus, our results will stimulate the study of new physical systems with cascade arrests leading to intermediate-scale condensates.
@@ Acknowledgements.
This work was supported in part by the French government through the UCAJEDI Investments in the Future project managed by the National Research Agency (ANR), with reference number ANR-15-IDEX-01, and by the International Centre for Theoretical Sciences (ICTS) for the online program - Turbulence: Problems at the Interface of Mathematics and Physics (code: ICTS/TPIMP2020/12). KVK thanks Laboratoire J. A. Dieudonné of Université Côte d’Azur for hospitality. RP and KVK acknowledge support from SERB and NSM (India) and thank SERC (IISc) for computational resources. KVK and DV are thankful to Siddhartha Mukherjee for useful discussions. DV acknowledges support from the International Centre for Theoretical Sciences (ICTS-TIFR), Bangalore, India.